# Composite likelihood estimation of stationary Gaussian processes

### v1

@@ Appendix A Mathematical proofs
We conclude from this and Lemma A.1 that $\text{var}(n^{-1/2}s_{n}(\theta))=O(L_{\gamma}(n))$, where $L_{\gamma}(n)=\sum_{l=1}^{n}L_{\infty}^{2}(l)l^{-1}$. By formula 1.5.8 and Proposition 1.5.9a in Bingham, Goldie, and Teugels (1989), $L_{\gamma}$ is slowly varying and growing faster than $L_{\infty}^{2}$, so the correct scaling of $s_{n}(\theta)$ is $n^{-1/2}L_{\gamma}^{-1/2}(n)$.
Next, we deduce the limiting distribution of $n^{-1/2}L_{\gamma}^{-1/2}(n)s_{n}(\theta)$. Following Hosking (1996) and Rosenblatt (1979), the $r$th order cross-cumulant of
$C_{k}=\sqrt{n}L_{\gamma}^{-1/2}(n)\left(\frac{1}{n}\sum_{t=1}^{n-k}Y_{t}Y_{t+k}\right),$
for $k=0,\dots,l$, is given by
$u_{r}=\frac{1}{2}\sum(r-1)!(nL_{\gamma}(n))^{-r/2}\sum_{j_{1}=1}^{n}\dots\sum_{j_{r}=1}^{n}\gamma_{j_{1}\pm\alpha_{1}-j_{2}}\gamma_{j_{2}\pm\alpha_{2}-j_{3}}\cdots\gamma_{j_{r-1}\pm\alpha_{r-1}-j_{r}}\gamma_{j_{r}\pm\alpha_{r}-j_{1}},$
where $\alpha_{i}\in\{0,1,\dots,l\}$, and the outer sum goes over all possible $2^{r}$ combinations of $+$ and -- signs in the equation above.
We shall show that for $r\geq 3$, $C_{r}\rightarrow 0$ as $n\rightarrow\infty$. As in Hosking (1996), we set $\alpha_{i}=0$ for convenience. Then, by Cauchy-Schwarz’s inequality
$|u_{r}|\leq 2^{r-1}(r-1)!(nL_{\gamma}(n))^{-r/2}\sum_{j_{1}=1}^{n}\dots\sum_{j_{r}=1}^{n}\gamma_{j_{1}-j_{2}}^{2}|\gamma_{j_{2}-j_{3}}\cdots|\gamma_{j_{r-1}-j_{r}}|.$
Now, for $j=1,\dots,n$,
$\begin{split}\sum_{i=1}^{n}|\gamma_{i-j}|&=\gamma_{0}+\sum_{i=1}^{j-1}|\gamma_{i}|+\sum_{i=1}^{n-j}|\gamma_{i}|\\ &\leq C\big{(}1+L_{\infty}(j)j^{1/2}+L_{\infty}(n-j)(n-j)^{1/2}\big{)}\\ &=O(L_{\infty}(n)n^{1/2}),\end{split}$
where Karamata’s integral inequality is used in the first passage. Moreover,
$\sum_{i=1}^{n}\gamma_{i-j}^{2}$ $=\gamma_{0}^{2}+\sum_{i=1}^{j-1}\gamma_{i}^{2}+\sum_{i=1}^{n-j}\gamma_{i}^{2}$
$\leq C\big{(}1+L_{\gamma}(j)+L_{\gamma}(n-j)\big{)}$
$=O(L_{\gamma}(n)).$
Holding these results together,
$u_{r}$ $=(nL_{\gamma}(n))^{-r/2}O(L_{\gamma}(n))O(n^{r/2-1}[L_{\infty}^{2}(n)]^{r/2-1})n$
$=O\big{(}L^{1-r/2}_{\gamma}(n)[L_{\infty}^{2}(n)]^{r/2-1}\big{)},$
following the remark above where $L_{\gamma}$ was introduced. So for $r\geq 3$, $u_{r}\rightarrow 0$ as $n\rightarrow\infty$. This implies a Gaussian limit, as claimed.
@@
$-\frac{1}{n}\frac{\partial}{\partial\theta^{\top}}s_{n}(\bar{\theta})\overset{\mathbb{P}}{\longrightarrow}H(\theta_{0}).$
@@
[…]
@@
$X_{n}=\sum_{k=1}^{n}x_{k}\text{ converges }\Leftrightarrow\tilde{X}_{n}=\sum_{k=1}^{n}\frac{n-k}{n}x_{k}\text{ converges.}$.
@@
To show the $\Leftarrow$ part, assume that $\tilde{X}_{n}$ converges. Then it is bounded by some number $C$. Suppose that $X_{n}$ does not converge, which means that it is unbounded, so there is an $N_{1}$ such that $X_{N_{1}}>2C$. Moreover, for any $N_{2}>2N_{1}$, the first $N_{1}$ terms in $\tilde{X}_{N_{2}}$ are at least half as large as the corresponding ones in $X_{N_{2}}$, since $21-N_{1}/N_{2}>1/$ 2. This implies that $\tilde{X}_{N_{1}}>\frac{1}{2}X_{N_{1}}>C$, which is a contradiction.
@@ Appendix B Method of moments-based estimator
As in the composite likelihood framework, the estimator of the mean, $\mu$, is the sample average, $\hat{\mu}=n^{-1}\sum_{i=1}^{n}Y_{i\Delta}$.
@@
Appendix C Monte Carlo analysis with $N=1$
Presented without comment.
Table 6: Parameter estimation of the fOU process ($\mu$ known).
Parameter Value MCLE MME
$T=1{,}095$ $T=1{,}825$ $T=2{,}555$ $T=1{,}095$ $T=1{,}825$ $T=2{,}555$
@@ Appendix C Monte Carlo analysis with $N=1$
$\mu$ 0.000 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000)
$\kappa$ 0.070 0.1051 (0.0502) 0.1002 (0.0375) 0.0912 (0.0290) 0.0716 (0.0217) 0.0702 (0.0167) 0.0696 (0.0139)
$\nu$ 0.200 0.2142 (0.0265) 0.2099 (0.0181) 0.2077 (0.0268) 0.1987 (0.0095) 0.1987 (0.0073) 0.1986 (0.0061)
$\alpha$ 0.200 0.1925 (0.0840) 0.1945 (0.0695) 0.1967 (0.0606) 0.1939 (0.0382) 0.1947 (0.0297) 0.1951 (0.0249)
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There is $N=1$ observation per unit interval, corresponding to a single observation every day. The true value of the parameter vector appear to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table reports the Monte Carlo average value of each parameter estimate across simulations (standard deviation in parenthesis).
Table 7: Parameter estimation of the fOU process ($\mu$ estimated).
Parameter Value MCLE MME
$T=1{,}095$ $T=1{,}825$ $T=2{,}555$ $T=1{,}095$ $T=1{,}825$ $T=2{,}555$
@@
$\mu$ 0.000 0.0004 (0.0684) -0.0004 (0.0591) -0.0003 (0.0532) 0.0003 (0.0681) -0.0003 (0.0586) -0.0002 (0.0528)
$\kappa$ 0.070 0.0841 (0.0228) 0.0846 (0.0181) 0.0790 (0.0147) 0.0776 (0.0224) 0.0745 (0.0169) 0.0730 (0.0140)
$\nu$ 0.200 0.2083 (0.0090) 0.2063 (0.0072) 0.2053 (0.0062) 0.1987 (0.0095) 0.1987 (0.0073) 0.1986 (0.0061)
$\alpha$ 0.200 0.1385 (0.0415) 0.1518 (0.0365) 0.1601 (0.0315) 0.1939 (0.0382) 0.1947 (0.0297) 0.1951 (0.0249)
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There is $N=1$ observation per unit interval, corresponding to a single observation every day. The true value of the parameter vector appear to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table reports the Monte Carlo average value of each parameter estimate across simulations (standard deviation in parenthesis).
Table 8: Parameter estimation of the Cauchy class ($\mu$ known).
Parameter Value MCLE MME
$T=1{,}095$ $T=1{,}825$ $T=2{,}555$ $T=1{,}095$ $T=1{,}825$ $T=2{,}555$
@@
$\mu$ 0.000 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000) 0.0000 (0.0000)
$\beta$ 1.250 1.3878 (0.2226) 1.3783 (0.1696) 1.3506 (0.1410) 1.0545 (0.1144) 1.0484 (0.0922) 1.0455 (0.0804)
$\nu$ 0.200 0.1757 (0.0054) 0.1757 (0.0041) 0.1820 (0.0036) 0.1999 (0.0062) 0.1999 (0.0048) 0.2000 (0.0040)
$\alpha$ 0.200 0.1323 (0.0977) 0.1295 (0.0754) 0.1485 (0.0648) -0.1490 (0.0421) -0.1481 (0.0326) -0.1475 (0.0275)
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There is $N=1$ observation per unit interval, corresponding to a single observation every day. The true value of the parameter vector appear to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table reports the Monte Carlo average value of each parameter estimate across simulations (standard deviation in parenthesis).
Table 9: Parameter estimation of the Cauchy class ($\mu$ estimated).
Parameter Value MCLE MME
$T=1{,}095$ $T=1{,}825$ $T=2{,}555$ $T=1{,}095$ $T=1{,}825$ $T=2{,}555$

### v2

@@ Appendix A Mathematical proofs
We conclude from this and Lemma A.1 that $\text{var}(n^{-1/2}s_{n}(\theta))=O(L_{\gamma}(n))$, where $L_{\gamma}(n)=\sum_{l=1}^{n}L_{\infty}^{2}(l\Delta)(l\Delta)^{-1}$. By formula 1.5.8 and Proposition 1.5.9a in Bingham, Goldie, and Teugels (1989), $L_{\gamma}$ is slowly varying and growing faster than $L_{\infty}^{2}$, so the correct scaling of $s_{n}(\theta)$ is $n^{-1/2}L_{\gamma}^{-1/2}(n)$.
To derive the limiting distribution, we rely on Theorem 4 in Hosking (1996), which shows the rather striking result that for $\beta\leq 1/2$, $\hat{\gamma}_{\ell}-\hat{\gamma}_{\ell^{\prime}}\overset{\mathbb{P}}{\longrightarrow}0$ for $\ell\neq\ell^{\prime}$. Hence, the sample autocovariances have a common asymptotic distribution at any lag, which is Gaussian for $\beta=1/2$, and their multivariate normal distribution function is degenerate. This implies that the entries in the score also exhibit this property, as they are a weighted sum of the sample autocovariances. Consequently,
$n^{-1/2}L_{\gamma}^{-1/2}(n)(\hat{\theta}_{\text{MCLE}}-\theta_{0})\overset{d}{\longrightarrow}\sum_{k^{j}\in Q}\sum_{j_{1},j_{2}\in k^{j}}\left[\frac{\partial}{\partial\theta_{r}}\big(\Sigma_{k}^{-1}(\theta)\big)_{j_{1},j_{2}}|_{\theta=\theta_{0}}\right]_{r=1}^{p}N(0,1).$
@@
$\frac{1}{n}\frac{\partial}{\partial\theta^{\top}}s_{n}(\bar{\theta})\overset{\mathbb{P}}{\longrightarrow}H(\theta_{0}).$
@@
$n^{\beta-1}L_{2}^{-1/2}(n)s_{n}(\theta_{0})\overset{d}{\longrightarrow}\frac{1}{2}\sum_{k^{j}\in Q}\sum_{j_{1},j_{2}\in k^{j}}\left[\frac{\partial}{\partial\theta_{r}}\big(\Sigma_{k}^{-1}(\theta)\big)_{j_{1},j_{2}}|_{\theta=\theta_{0}}\right]_{r=1}^{p}Z_{2,H}(1).$
@@
$X_{n}=\sum_{k=1}^{n}x_{k}\text{ converges }\Leftrightarrow\tilde{X}_{n}=\sum_{k=1}^{n}\frac{n-k}{n}x_{k}\text{ converges.}$
@@
To show the $\Leftarrow$ part, assume that $\tilde{X}_{n}$ converges. Then it is bounded by some number $C$. Suppose that $X_{n}$ does not converge, which means that it is unbounded, so there is an $N_{1}$ such that $X_{N_{1}}>2C$. Moreover, for any $N_{2}>2N_{1}$, the first $N_{1}$ terms in $\tilde{X}_{N_{2}}$ are at least half as large as the corresponding ones in $X_{N_{2}}$, since $1-N_{1}/N_{2}>1/2$. This implies that $\tilde{X}_{N_{1}}>\frac{1}{2}X_{N_{1}}>C$, which is a contradiction.
@@ Appendix B Method of moments-based estimator
The estimator of the mean, $\mu$, is the sample average,
$\hat{\mu}=\frac{1}{n}\sum_{i=1}^{n}Y_{i\Delta}.$
@@
Appendix C Additional Monte Carlo analysis
In this appendix, we provide some supplemental simulation evidence to highlight the performance of our MCLE approach under alternative configurations:
  • Table 6 – Table 9 report the pairwise setting ($q=2$ and $N=12$).
  • Table 10 – Table 13 consider the setting with daily data ($q=3$ and $N=1$).
The tables are presented without comment.
Table 6: Parameter estimation of the fOU process [$q=2$ and $\mu$ known].
Parameter Value T=1,095T=1{,}095 T=1,825T=1{,}825 T=2, [^5] T=2{,}555
MCLE MME RMSEr MCLE MME RMSEr MCLE MME RMSEr
@@ Appendix C Additional Monte Carlo analysis
$\mu$ 0.000
$\kappa$ 0.070 0.005 (0.019) 0.003 (0.016) 1.202 0.003 (0.014) 0.002 (0.012) 1.183 0.002 (0.012) 0.001 (0.010) 1.182
$\nu$ 0.200 0.001 (0.008) 0.000 (0.008) 0.998 0.001 (0.006) -0.000 (0.006) 0.986 0.000 (0.005) -0.000 (0.005) 0.976
$\alpha$ 0.200 0.001 (0.015) -0.000 (0.011) 1.362 0.001 (0.011) -0.000 (0.008) 1.359 0.001 (0.010) -0.000 (0.007) 1.348
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval corresponding to twelve every day. The true value of the parameter vector appears to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table shows the Monte Carlo average of each parameter estimate across simulations (standard deviation in parenthesis). The last column reports the RMSE ratio defined as $\text{RMSE}_{r}=\text{RMSE}(\hat{\theta}_{\text{MCLE}})/\text{RMSE}(\hat{\theta}_{\text{MME}})$.
Table 7: Parameter estimation of the fOU process [$q=2$ and $\mu$ estimated].
Parameter Value T=1,095T=1{,}095 T=1,825T=1{,}825 T=2, [^5] T=2{,}555
MCLE MME RMSEr MCLE MME RMSEr MCLE MME RMSEr
@@
$\mu$ 0.000 0.000 (0.069) 0.000 (0.069) 1.000 0.000 (0.059) 0.000 (0.059) 1.000 0.000 (0.054) 0.000 (0.054) 1.000
$\kappa$ 0.070 0.013 (0.019) 0.009 (0.016) 1.238 0.009 (0.014) 0.006 (0.012) 1.213 0.007 (0.012) 0.005 (0.010) 1.210
$\nu$ 0.200 0.003 (0.008) 0.000 (0.008) 1.045 0.002 (0.006) -0.000 (0.006) 1.019 0.002 (0.005) -0.000 (0.005) 1.002
$\alpha$ 0.200 0.005 (0.015) -0.000 (0.011) 1.388 0.003 (0.011) -0.000 (0.008) 1.376 0.003 (0.009) -0.000 (0.007) 1.361
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval corresponding to twelve every day. The true value of the parameter vector appears to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table shows the Monte Carlo average of each parameter estimate across simulations (standard deviation in parenthesis). The last column reports the RMSE ratio defined as $\text{RMSE}_{r}=\text{RMSE}(\hat{\theta}_{\text{MCLE}})/\text{RMSE}(\hat{\theta}_{\text{MME}})$.
Table 8: Parameter estimation of the Cauchy class [$q=2$ and $\mu$ known].
Parameter Value T=1,095T=1{,}095 T=1,825T=1{,}825 T=2, [^5] T=2{,}555
MCLE MME RMSEr MCLE MME RMSEr MCLE MME RMSEr
@@
$\mu$ 0.000
$\beta$ 1.250 0.010 (0.113) 0.040 (0.184) 0.606 0.006 (0.087) 0.030 (0.146) 0.592 0.004 (0.074) 0.026 (0.125) 0.586
$\nu$ 0.200 -0.000 (0.006) -0.000 (0.006) 1.002 -0.000 (0.005) -0.000 (0.005) 1.002 -0.000 (0.004) -0.000 (0.004) 1.002
$\alpha$ 0.200 0.000 (0.010) 0.041 (0.011) 0.326 0.000 (0.008) 0.041 (0.008) 0.259 0.000 (0.007) 0.041 (0.007) 0.222
Note. We simulate the process in the caption of the table 10,000 times on the interval $[0,T]$, where $T$ is interpreted as the number of days. There are $N=12$ observations per unit interval corresponding to twelve every day. The true value of the parameter vector appears to the left in Panel A – E. We estimate $\theta$ with the maximum composite likelihod estimation (MCLE) procedure developed in the main text, and benchmark it against a method-of-moments estimator (MME). The table shows the Monte Carlo average of each parameter estimate across simulations (standard deviation in parenthesis). The last column reports the RMSE ratio defined as $\text{RMSE}_{r}=\text{RMSE}(\hat{\theta}_{\text{MCLE}})/\text{RMSE}(\hat{\theta}_{\text{MME}})$.
Table 9: Parameter estimation of the Cauchy class [$q=2$ and $\mu$ estimated].
Parameter Value T=1,095T=1{,}095 T=1,825T=1{,}825 T=2, [^5] T=2{,}555
MCLE MME RMSEr MCLE MME RMSEr MCLE MME RMSEr